# A Doppler backscattering diagnostic for the EXL-50U spherical tokamak: plasma considerations and preliminary quasioptical design

### v3

@@ 5.2 Spatial resolution
6 Further improvements
In this section, we briefly discuss how the quasioptics can be further improved.
An alternative design uses two in-vessel mirrors to bring the launch position closer to the plasma and nearer to the midplane, allow better access to the plasma, see Fig. 17. This design would result in the beam probing the cutoff at lower angles, allowing lower wavenumbers to be measured. However, this requires detailed consideration of in-vessel space constraints and is beyond the scope of this work.
Figure 17: A possible two-mirror DBS design that can be implemented for the EXL-50U. These mirrors can be adjusted to ensure the beams are not incident on PF14, providing easier access to the plasma and increasing the maximum poloidal launch angle range. The mirrors are concave, providing a focusing effect for the beams.
As the edge pitch angle of spherical tokamaks can vary significantly with operation, we may also add a polariser to fine-tune coupling of the probe beam as it propagates from the vacuum into the plasma.
@@ 7 Conclusion
We completed a quasioptical design for a DBS for the EXL-50U spherical tokamak. Using SCOTTY simulations, we show that the DBS can access turbulent fluctuations in H-mode (A) over a wide range of scattering locations, $0.30<\rho<1$, thereby providing access to the pedestal, the ITB region, and parts of the core. It is also able to measure $0.24\textrm{ mm}^{-1}<k_{\perp}<0.95\textrm{ mm}^{-1}$. We also show that poloidal flux coverage of the H-mode (B) and L-mode profiles is adequate. Across all profiles, the DBS can measure turbulence on the scale of $k_{\perp}\rho_{s}$ $\lesssim$ 10 with a maximum range of $0.72<k_{\perp}\rho_{s}<10.50$. This makes it suitable for measuring ion-scale turbulence and the lower end of electron-scale turbulence (10 $\lesssim$ $k_{\perp}\rho_{s}$ $\lesssim$ 30), which is relevant to EXL-50U as it is a spherical tokamak. To probe these turbulent fluctuations, the DBS must operate in the U-band frequency range and a poloidal launch angle range of $43.5^{\circ} \leq$ $\varphi_{p}$ $\leq$ $35.5^{\circ}$. In this work, a quasioptical system operating at 40–60 GHz was designed while satisfying physical constraints.
@@ 8 Acknowledgements
This research was supported by A*STAR: by the FEAT SRTT and a SERC Central Research Fund. Y. H. M. Liang was funded by a National Science Scholarship from A*STAR, Singapore. This material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, under Award DE-SC0021150.

### v4

@@ 6 Conclusion
We completed a quasioptical design for a DBS for the EXL-50U spherical tokamak. Using SCOTTY simulations, we show that the DBS can access turbulent fluctuations in H-mode (A) over a wide range of scattering locations, $0.30<\rho<1$, thereby providing access to the pedestal, the ITB region, and parts of the core. It is also able to measure $0.24\textrm{ mm}^{-1}<k_{\perp}<0.95\textrm{ mm}^{-1}$. We also show that poloidal flux coverage of the H-mode (B) and L-mode profiles is adequate. Across all profiles, the DBS can measure turbulence on the scale of $k_{\perp}\rho_{s}$ $\lesssim$ 10 with a maximum range of $0.72<k_{\perp}\rho_{s}<10.50$. This makes it suitable for measuring ion-scale turbulence and, in certain cases, the lower end of electron-scale turbulence (10 $\lesssim$ $k_{\perp}\rho_{s}$ $\lesssim$ 30), which might potentially enable investigation of cross-scale turbulent transport, which is expected to be important in spherical tokamaks. To probe these turbulent fluctuations, the DBS must operate in the U-band frequency range and a poloidal launch angle range of $43.5^{\circ} \leq$ $\varphi_{p}$ $\leq$ $35.5^{\circ}$. In this work, a quasioptical system operating at 40–60 GHz was designed while satisfying physical constraints. The present design targets mainly ion-scale turbulence. To measure the higher end of electron-scale turbulence and fully understand cross-scale interactions, further optimisation of the quasioptical system is required.
@@ 7 Acknowledgements
This research was supported by A*STAR through the FEAT SRTT and the SERC Central Research Fund, and was partially funded by the ENN group. Ying Hao Matthew Liang was funded by a National Science Scholarship from A*STAR, Singapore. The material is also based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, under Award DE-SC0021150. This work is also supported by the High-End Talents Program of Hebei Province , Innovative Approaches towards Development of Carbon Free Clean Fusion Energy (No. 2021HBQZYCSB006).